OpenStudy (osanseviero):

Proof by induction of this:

OpenStudy (osanseviero):

cosx + cos2x + ... + cosnx =\[\frac{ \cos[\frac{ x }{ 2 }(n+1)]sen \frac{ nx }{ 2 }}{ sen(\frac{ x }{ 2} }\]

OpenStudy (osanseviero):

I did base caseand some induction, but it seems that my algebra is not helping

OpenStudy (freckles):

what is sen?

OpenStudy (anonymous):

original sin

OpenStudy (freckles):

secant?

OpenStudy (freckles):

oh sine

OpenStudy (osanseviero):

sorry, sine

OpenStudy (freckles):

so is the following true: the case for n=1 \[\frac{\cos(\frac{x}{2}(1+1))\sin(\frac{1x}{2})}{\sin(\frac{x}{2})}=\cos(x)\] base case is always the easiest The induction part might actually be hard I haven't tried yet

OpenStudy (anonymous):

maybe when you add you can use some trig identity right?

OpenStudy (anonymous):

oh wait i screwed that up

OpenStudy (anonymous):

\[\frac{\cos(\frac{x}{2}(k+1))\sin(\frac{kx}{2})}{\sin(\frac{x}{2})}+\cos((k+1)x)=\frac{\cos(\frac{x}{2}(k+2))\sin(\frac{(k+1)x}{2})}{\sin(\frac{x}{2})}\]

OpenStudy (anonymous):

some identity with \(\sin(\frac{x}{2})\cos((k+1)x)\)

OpenStudy (anonymous):

it is half of something \[\sin(\alpha)\cos(\beta)=\frac{1}{2}\sin(\alpha+\beta)\] maybe that will do it

OpenStudy (anonymous):

2 sin a cos b=sin (a+b)+sin (a-b)

OpenStudy (osanseviero):

I've tried that already, no luck

OpenStudy (osanseviero):

I will try again